OpenStudy (anonymous):

If an associate’s average sale to 5 customers is $125, and her sales to four of the customers are $200, $150, $125, and $75, what is the associate’s sale to the fifth customer?

OpenStudy (anonymous):

$75

OpenStudy (anonymous):

125 * 5 = 625 200 + 150 + 125 + 75 = 550 625 - 550 = 75 $75

OpenStudy (stormfire1):

Since the average is the sum of all the values divided by the total number of values, you can simply put it in the equation and solve for the missing value: \[(200+150+125+75+x)/5 = 125\]\[625 = 200+150+125+75+x\]\[625=550+x\]\[x=75\]